OpenStudy (anonymous):

what is radical 24 with index 4...?

OpenStudy (anonymous):

index 3 my bad

jimthompson5910 (jim_thompson5910):

Hint: the perfect cubes are 1, 8, 27, 64, etc..

OpenStudy (anonymous):

o thanks

jimthompson5910 (jim_thompson5910):

do you see what to do?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

im trying to figure it out but idk

jimthompson5910 (jim_thompson5910):

factor 24 into 8*3 then use the rule \[\Large \sqrt[3]{xy} = \sqrt[3]{x}*\sqrt[3]{y}\]

OpenStudy (anonymous):

why 8?

jimthompson5910 (jim_thompson5910):

because 8 is a perfect cube (since \(\Large 2^3 = 8\) )

OpenStudy (anonymous):

oh okay thanks

jimthompson5910 (jim_thompson5910):

8 is the largest factor of 24 (that's also a perfect cube)

OpenStudy (anonymous):

oh alrightt :)